OpenStudy (anonymous):

Determine this difference. 8.54 - (-3.76)

OpenStudy (alexandervonhumboldt2):

-(-)=+

OpenStudy (alexandervonhumboldt2):

8.54-(-3.76)=8.54+3.76=?

OpenStudy (anonymous):

I'm bad at this, and i need you to just tell me plz

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

tell what? you can't add 8.54+3.76?

OpenStudy (anonymous):

sorry i didn't understand